27.27 A 10.0-uF capacitor is charged by a 10.0-V battery through a resistance R. The capacitor reaches a potential difference of 4.00 V in a time interval of 3.00 s after charging begins. Find R.
1
Expert's answer
2020-04-13T10:15:51-0400

The potential difference on the capacitor

"V(t)=V_s\left(1-\exp(-t/RC)\right)."

Hence, the resistance

"R=-\frac{t}{C\ln(1-V/V_s)}=""=-\frac{3.00}{10.0\times 10^{-6}\ln(1-4.00/10.0)}=""=1.96\times 10^6\: \Omega=1.96\:\rm M\Omega"